how do you find the roots of 3x^2-5x=2 ? and could you please show me how to work out the problem?
Answer by scott8148(6628)

3x^2 - 5x - 2 = 0

factoring ___ (3x + 1)(x - 2) = 0

3x + 1 = 0 ___ x = -1/3

x - 2 = 0 ___ x = 2